OpenStudy (kainui):

Why is sine and cosine defined at 90 degrees? It makes a 90-90-0 triangle which makes no sense tome.

OpenStudy (kainui):



OpenStudy (experimentx):

why wouldn't it defined?

OpenStudy (kainui):

For example, in my mind sine of 90 degrees would be \[\frac{ \infty }{ \infty }\] because the hypotenuse would be parallel to the opposite side of the triangle separated by 0 distance. I guess we can say they intersect since they're the same line and dividing infinity by itself =1 but the whole thing doesn't sit well with me... I mean, should I really accept these premises as to why sin(90)=1? Do I not really understand what the sine function does, and the triangle idea is just a way to explain it to people learning?

OpenStudy (experimentx):

why ... does your hypotenuse have infinite length?

OpenStudy (experimentx):